# Open dynamical systems with a moving hole

### v1

@@ 3. Dimensions of $K^{\omega}$ for progressively overlapping or totally distinct $\omega$
In this section we will prove the dimension equality of $K^{\omega}$ for progressively overlapping and totally distinct sequences $\omega\in(D_{b}^{m})^{\mathbb{N}}$, and establish Theorem 1.4. Recall the definitions of progressively overlapping and totally distinct sequences in subsection 1.1, and keep in mind that $b\in\mathbb{N}_{\geq 3}$ and $m\in\mathbb{N}_{\geq 2}$.
@@
For $\omega\in(D_{b}^{m})^{\mathbb{N}}$, if there exist $\beta>1$ and $C>1$ such that
@@
We first show the upper bound in (3.1). Since $r<b^{-k_{0}}$, the ball $B(x,r)$ intersects at most three intervals in $\{I_{\mathbf{d}}:\mathbf{d}\in\Sigma_{k_{0}}^{\omega}\}$, denoted by $I_{\mathbf{c}_{1}},I_{\mathbf{c}_{2}},I_{\mathbf{c}_{3}}$ with $\mathbf{c}_{i}\in\Sigma_{k_{0}}^{\omega}$ for $1\leq i\leq 3$. That is, $K^{\omega}\cap B(x,r)\subset I_{\mathbf{c}_{1}}\cup I_{\mathbf{c}_{2}}\cup I_{\mathbf{c}_{3}}$. Recall the definition of $\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{i})$ in (2.2). For $n\in\mathbb{N}$, the family of intervals $\mathcal{I}_{n}:=\{I_{\mathbf{d}}:\mathbf{d}\in\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{1})\cup\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{2})\cup\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{3})\}$ forms a $b^{-(k_{0}+n)}$-covering of $K^{\omega}\cap B(x,r)$. By Lemma 2.2 and (3.2), we have
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}}$ and any $k\in\mathbb{N}$ we have
@@
Write $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots\in(D_{b}^{m})^{\mathbb{N}}$ with each $\omega^{i}=\omega_{1}^{i}\omega_{2}^{i}\ldots\omega_{m}^{i}\in D_{b}^{m}$, and fix $k\in\mathbb{N}$.
@@
(ii) In this part, we prove the upper bound. Note by (2.1) that each $d_{1}d_{2}\ldots d_{k+m}\in\Sigma_{k+m}^{\omega}$ satisfies $d_{k+1}d_{k+2}\ldots d_{k+m}\neq\omega_{1}^{k}\omega_{2}^{k}\ldots\omega_{m}^{k}$. Conditioned on the maximum index $j$ such that $d_{k+j}\neq\omega_{j}^{k}$ we make the following partition of $\Sigma_{k+m}^{\omega}$ as
@@
$x^{m}=(b-1)(x^{m-1}+x^{m-2}+\ldots+x+1)\quad\textrm{and}\quad x^{m}=bx^{m-1}-1,$
@@
Let $f(x):=x^{m}-(b-1)(x^{m-1}+x^{m-2}+\ldots+x+1)$. It is easy to verify that $f(b-1)<0$ and $f(x)\geq f(b)=1$ for all $x\geq b$. By the intermediate value theorem the largest real root $\lambda\in(b-1,b)$. Write $\widetilde{f}(x):=(x-1)f(x)=x^{m+1}-bx^{m}+b-1=f_{1}(x)-f_{2}(x)$ with $f_{1}(x):=x^{m+1}$ and $f_{2}(x):=bx^{m}-(b-1)$. Note that
@@
which implies that $z=1$. Note by using $m\in\mathbb{N}_{\geq 2},b\in\mathbb{N}_{\geq 3}$ that the derivative $\widetilde{f}^{\prime}(1)\neq 0$. Thus, $z=1$ is the unique zero of $\widetilde{f}$ on the unit circle. This implies that $f(x)=\widetilde{f}(x)/(x-1)$ has $m-1$ zeros inside the unit circle. It follows that $\lambda$ is a Pisot number.
Let $g(x):=x^{m}-bx^{m-1}+1$. It is easy to check $g(b-1)<0$ and $g(x)\geq g(b)=1$ for all $x\geq b$. By the intermediate value theorem the largest real root $\eta\in(b-1,b)$. Write $g_{1}(x):=bx^{m-1}-1$ and $g_{2}(x):=x^{m}$. Note by $b\geq 3$ that
$|g_{1}(z)|=|bz^{m-1}-1|\geq b-1>1=|g_{2}(z)|\quad\text{for any}\leavevmode\nobreak\ z\in\mathbb{C}\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ |z|=1.$
By Rouche’s theorem, $g_{1}(x)$ and $g(x)=g_{2}(x)-g_{1}(x)$ have the same number of zeros inside the unit circle. Since $g_{1}(x)=bx^{m-1}-1$ has $m-1$ zeros inside the unit circle, so is $g(x)$. It follows that $\eta$ is a Pisot number. Since $1<\lambda<b$ and $\lambda^{m}(\lambda-1)=(b-1)(\lambda^{m}-1)$, we have
@@
which implies that $g(\lambda)>0$. By the definition of $\eta$ we conclude that $\eta<\lambda$, completing the proof. ∎
For $\mathbf{v}=(v_{1},v_{2},\ldots,v_{n})\in\mathbb{R}^{n}\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \mathbb{C}^{n}$, let $\|\mathbf{v}\|:=\sum_{i=1}^{n}|v_{i}|$. Set
(3.9) $V:=\big{\{}\mathbf{v}=(v_{m},\ldots,v_{1})\in\mathbb{R}^{m}:\|\mathbf{v}\|=1,\leavevmode\nobreak\ v_{1}>0,\leavevmode\nobreak\ (b-1)v_{i}\leq v_{i+1}\leq bv_{i}\leavevmode\nobreak\ \leavevmode\nobreak\ \forall 1\leq i<m\big{\}}.$
@@
It’s easy to check that the characteristic polynomials of $A$ and $B$ are $f(x)=x^{m}-(b-1)(x^{m-1}+x^{m-2}+\ldots+x+1)$ and $g(x)=x^{m}-bx^{m-1}+1$, respectively. So, by Lemma 3.3 it follows that $\lambda$ and $\eta$ are the spectral radii of $A$ and $B$, respectively.
@@
where $\lambda$ is the largest real root of $x^{m}=(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)$. Note that for any $\mathbf{v}=(v_{m},\ldots,v_{1})\in V$ we have $\delta_{0}\leq v_{i}<1$ for all $1\leq i\leq m,$ where $\delta_{0}=(1+b+\ldots+b^{m-1})^{-1}>0$. Therefore, we conclude that for $n\in\mathbb{N}$ and $\mathbf{v}\in V$,
@@
Next, we focus on the matrix $B$. We first show that for any $\mathbf{v}\in V$ the sequence $\{\|\mathbf{v}B^{n}\|\}_{n=0}^{\infty}$ is increasing. Write $\mathbf{v}=(v_{m},\ldots,v_{1})\in V$. For $n\in\mathbb{N}$, we recursively define $v_{n+m}=bv_{n+m-1}-v_{n}$. Then we have $\mathbf{v}B^{n}=(v_{n+m},\ldots,v_{n+1})$ for all $n\in\mathbb{N}_{0}$. Note that $v_{m}>v_{m-1}>\ldots>v_{1}$. Then $v_{m+1}=bv_{m}-v_{1}=(b-1)v_{m}+(v_{m}-v_{1})>v_{m}$. By induction, we can show that the sequence $\{v_{n}\}_{n=1}^{\infty}$ is increasing. It follows that the sequence $\{\|\mathbf{v}B^{n}\|\}_{n=0}^{\infty}$ is increasing.
Write $g(x)=x^{m}-bx^{m-1}+1$ for the characteristic polynomial of $B$. Note that all roots of equation $g^{\prime}(x)=0$ are $x=0$ and $x=b(m-1)/m>1$. By Lemma 3.3, $x=\eta$ is the unique root of equation $g(x)=0$ outside the unit circle. Thus, the equations $g(x)=0$ and $g^{\prime}(x)=0$ do not share a common root. This means that the equation $g(x)=0$ has $m$ distinct roots in $\mathbb{C}$, denoted by $\eta,\eta_{1},\ldots,\eta_{m-1}$. By Lemma 3.3, we have $|\eta_{i}|<1$ for all $11\leq i\leq m-$ 1. Let $\mathbf{v}^{0},\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{m-1}\in\mathbb{C}^{m}$ with $\|\mathbf{v}_{i}\|=1$ be the left eigenvectors corresponding to $\eta,\eta_{1},\eta_{2},\ldots,\eta_{m-1}$, respectively. That is, $\mathbf{v}^{0}B=\eta\mathbf{v}^{0}$ and $\mathbf{v}^{i}B=\eta_{i}\mathbf{v}^{i}$ for all $11\leq i\leq m-$ 1. Then $\mathbf{v}^{0},\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{m-1}$ are linearly independent in $\mathbb{C}^{m}$. So, each $\mathbf{v}\in V$ can be uniquely written as
@@
Note that $|\eta_{i}|<1$ for all $11\leq i\leq m-$ 1. We obtain that
@@
Note that $V$ is a compact set in $\mathbb{C}^{m}$, and for $10\leq i\leq m-$ 1 the coefficient $c_{i}(\mathbf{v})$ is continuous as a function of $\mathbf{v}$. It follows from the claim that
@@
[…]
@@
Therefore, by taking $C_{1}=\max\{\eta^{n_{0}},3\widetilde{c}_{0}\eta^{n_{0}}/2,2/c_{0},3\widetilde{c}_{0}/2,c/(m^{2}\delta_{0}),cm^{2}\}$, we complete the proof. ∎
@@
There exists a constant $C_{2}>1$ such that for any progressively overlapping sequence $\zeta\in(D_{b}^{m})^{\mathbb{N}}$ and for any totally distinct sequence $\xi\in(D_{b}^{m})^{\mathbb{N}}$, we have
$C_{2}^{-1}\lambda^{k}\leq|\Sigma_{k}^{\zeta}|\leq C_{2}\lambda^{k}\quad\text{and}\quad C_{2}^{-1}\eta^{k}\leq|\Sigma_{k}^{\xi}|\leq C_{2}\eta^{k}\quad\forall k\in\mathbb{N}.$
@@
where $\widetilde{\mathbf{v}}_{0}=(|\Sigma_{m}^{\xi}|,|\Sigma_{m-1}^{\xi}|,\ldots,|\Sigma_{1}^{\xi}|)$
@@
For $n\in\mathbb{N}$, let $\zeta\in(D_{b}^{m})^{\mathbb{N}}$ be progressively overlapping at position-$k$ for all $1\leq k\leq n$, and let $\xi\in(D_{b}^{m})^{\mathbb{N}}$ be totally distinct at position-$k$ for all $1\leq k\leq n$. Then by the proof of Proposition 3.5, there exists a constant $C_{2}>1$ such that

### v2

@@ 3. Dimensions of $K^{\omega}$ for progressively overlapping or totally distinct $\omega$
In this section we will prove the dimension equality of $K^{\omega}$ for progressively overlapping and totally distinct sequences $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, and establish Theorem 1.4. Recall the definitions of progressively overlapping and totally distinct sequences in subsection 1.1, and keep in mind that $b\in\mathbb{N}_{\geq 3}$ and $m\in\mathbb{N}_{\geq 2}$.
@@
For $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$, if there exist $\beta>1$ and $C>1$ such that
@@
We first show the upper bound in (3.1). Since $r<b^{-k_{0}}$, the ball $B(x,r)$ intersects at most three intervals in $\{I_{\mathbf{d}}:\mathbf{d}\in\Sigma_{k_{0}}^{\omega}\}$, denoted by $I_{\mathbf{c}_{1}},I_{\mathbf{c}_{2}},I_{\mathbf{c}_{3}}$ with $\mathbf{c}_{i}\in\Sigma_{k_{0}}^{\omega}$ for $1\leq i\leq 3$. That is, $K^{\omega}\cap B(x,r)\subset I_{\mathbf{c}_{1}}\cup I_{\mathbf{c}_{2}}\cup I_{\mathbf{c}_{3}}$. Recall the definition of $\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{i})$ in (2.2). For $n\in\mathbb{N}$, the family of intervals $\mathcal{I}_{n}:=\{I_{\mathbf{d}}:\mathbf{d}\in\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{1})\cup\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{2})\cup\Sigma_{k_{0}+n}^{\omega}(\mathbf{c}_{3})\}$ forms a $b^{-(k_{0}+n)}$-covering of $K^{\omega}\cap B(x,r)$. By the assumption and using Lemma 2.2 and (3.2), we have
@@
For any $\omega\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and any $k\in\mathbb{N}$ we have
@@
Write $\omega=\omega^{0}\omega^{1}\omega^{2}\ldots\in(D_{b}^{m})^{\mathbb{N}_{0}}$ with each $\omega^{i}=\omega_{1}^{i}\omega_{2}^{i}\ldots\omega_{m}^{i}\in D_{b}^{m}$, and fix $k\in\mathbb{N}$.
@@
(ii) Now we prove the upper bound. Note by (2.1) that each $d_{1}d_{2}\ldots d_{k+m}\in\Sigma_{k+m}^{\omega}$ satisfies $d_{k+1}d_{k+2}\ldots d_{k+m}\neq\omega_{1}^{k}\omega_{2}^{k}\ldots\omega_{m}^{k}$. Conditioned on the maximum index $j$ such that $d_{k+j}\neq\omega_{j}^{k}$ we make the following partition of $\Sigma_{k+m}^{\omega}$ as
@@
$x^{m}=(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)\quad\textrm{and}\quad x^{m}=bx^{m-1}-1,$
@@
Let $f(x):=x^{m}-(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)$. It is easy to verify that f(b1)<0f(b-1)<0 and $f(x)\geq f(b)=1$ for all $x\geq b$. By the intermediate value theorem the largest real root $\lambda\in(b-1,b)$. Write $\widetilde{f}(x):=(x-1)f(x)=x^{m+1}-bx^{m}+b-1=f_{1}(x)-f_{2}(x)$ with $f_{1}(x):=x^{m+1}$ and $f_{2}(x):=bx^{m}-(b-1)$. Note that
@@
which implies that $z=1$. Note by using $m\in\mathbb{N}_{\geq 2},b\in\mathbb{N}_{\geq 3}$ that the derivative $\widetilde{f}^{\prime}(1)\neq 0$. Thus, $z=1$ is the unique zero of $\widetilde{f}$ on the unit circle. This implies that $f(x)=\widetilde{f}(x)/(x-1)$ has m1m-1 zeros inside the unit circle. It follows that $\lambda$ is a Pisot number.
Let $g(x):=x^{m}-bx^{m-1}+1$. It is easy to check g(b1)<0g(b-1)<0 and $g(x)\geq g(b)=1$ for all $x\geq b$. By the intermediate value theorem the largest real root $\eta\in(b-1,b)$. Write $g_{1}(x):=bx^{m-1}-1$ and $g_{2}(x):=x^{m}$. Note by $b\geq 3$ that
$|g_{1}(z)|=|bz^{m-1}-1|\geq b-1>1=|g_{2}(z)|\quad\text{for any}\penalty 10000\ z\in\mathbb{C}\penalty 10000\ \text{with}\penalty 10000\ |z|=1.$
By Rouche’s theorem, $g_{1}(x)$ and $g(x)=g_{2}(x)-g_{1}(x)$ have the same number of zeros inside the unit circle. Since $g_{1}(x)=bx^{m-1}-1$ has m1m-1 zeros inside the unit circle, so is $g(x)$. It follows that $\eta$ is a Pisot number. Since $1<\lambda<b$ and $\lambda^{m}(\lambda-1)=(b-1)(\lambda^{m}-1)$, we have
@@
which implies that $g(\lambda)>0$. By the definition of $\eta$ we conclude that $\eta<\lambda$. ∎
For $\mathbf{v}=(v_{1},v_{2},\ldots,v_{n})\in\mathbb{R}^{n}\penalty 10000\ \text{or}\penalty 10000\ \mathbb{C}^{n}$, let $\|\mathbf{v}\|:=\sum_{i=1}^{n}|v_{i}|$. Set
(3.9) $V:=\big\{\mathbf{v}=(v_{m},\ldots,v_{1})\in\mathbb{R}^{m}:\|\mathbf{v}\|=1,\penalty 10000\ v_{1}>0,\penalty 10000\ (b-1)v_{i}\leq v_{i+1}\leq bv_{i}\penalty 10000\ \penalty 10000\ \forall 1\leq i<m\big\}.$
@@
It’s easy to check that the characteristic polynomials of $A$ and $B$ are $f(x)=x^{m}-(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)$ and $g(x)=x^{m}-bx^{m-1}+1$, respectively. So, by Lemma 3.3 it follows that $\lambda$ and $\eta$ are the spectral radii of $A$ and $B$, respectively.
@@
where $\lambda$ is the largest real root of $x^{m}=(b-1)(x^{m-1}+x^{m-2}+\cdots+x+1)$. Note that for any $\mathbf{v}=(v_{m},\ldots,v_{1})\in V$ we have $\delta_{0}\leq v_{i}<1$ for all $1\leq i\leq m,$ where $\delta_{0}=(1+b+\cdots+b^{m-1})^{-1}>0$. Therefore, we conclude that for $n\in\mathbb{N}$ and $\mathbf{v}\in V$,
@@
Next, we focus on the matrix $B$. We first show that for any $\mathbf{v}\in V$ the sequence $\{\|\mathbf{v}B^{n}\|\}_{n=0}^{\infty}$ is increasing. Write $\mathbf{v}=(v_{m},\ldots,v_{1})\in V$. For $n\in\mathbb{N}$, we recursively define $v_{n+m}=bv_{n+m-1}-v_{n}$. Then we have $\mathbf{v}B^{n}=(v_{n+m},\ldots,v_{n+1})$ for all $n\in\mathbb{N}_{0}$. Note that $v_{m}>v_{m-1}>\cdots>v_{1}$. Then $v_{m+1}=bv_{m}-v_{1}=(b-1)v_{m}+(v_{m}-v_{1})>v_{m}$. By induction, we can show that the sequence $\{v_{n}\}_{n=1}^{\infty}$ is increasing. It follows that the sequence $\{\|\mathbf{v}B^{n}\|\}_{n=0}^{\infty}$ is increasing.
Write $g(x)=x^{m}-bx^{m-1}+1$ for the characteristic polynomial of $B$. Note that all roots of equation $g^{\prime}(x)=0$ are $x=0$ and x=b(m1)/m>1x=b(m-1)/m>1. By Lemma 3.3, $x=\eta$ is the unique root of equation $g(x)=0$ outside the unit circle. Thus, the equations $g(x)=0$ and $g^{\prime}(x)=0$ do not share a common root. This means that the equation $g(x)=0$ has $m$ distinct roots in $\mathbb{C}$, denoted by $\eta,\eta_{1},\ldots,\eta_{m-1}$. By Lemma 3.3, we have $|\eta_{i}|<1$ for all $1\leq i\leq m-1$. Let $\mathbf{v}^{0},\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{m-1}\in\mathbb{C}^{m}$ with $\|\mathbf{v}_{i}\|=1$ be the left eigenvectors corresponding to $\eta,\eta_{1},\eta_{2},\ldots,\eta_{m-1}$, respectively. That is, $\mathbf{v}^{0}B=\eta\mathbf{v}^{0}$ and $\mathbf{v}^{i}B=\eta_{i}\mathbf{v}^{i}$ for all $1\leq i\leq m-1$. Then $\mathbf{v}^{0},\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{m-1}$ are linearly independent in $\mathbb{C}^{m}$. So, each $\mathbf{v}\in V$ can be uniquely written as
@@
Note that $|\eta_{i}|<1$ for all $1\leq i\leq m-1$. We obtain that
@@
Note that $V$ is a compact set in $\mathbb{C}^{m}$, and for $0\leq i\leq m-1$ the coefficient $c_{i}(\mathbf{v})$ is continuous as a function of $\mathbf{v}$. It follows from the claim that
@@
Note that $\sup\{|c_{i}(\mathbf{v})|:\mathbf{v}\in V,\penalty 10000\ 1\leq i\leq m-1\}<\infty$. For $\mathbf{v}\in V$, we have $\mathbf{v}B^{n}=c_{0}(\mathbf{v})\eta^{n}\mathbf{v}^{0}+\sum_{i=1}^{m-1}c_{i}(\mathbf{v})\eta_{i}^{n}\mathbf{v}^{i}$. Therefore, there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and for all $\mathbf{v}\in V$,
@@
Hence, by taking $C_{1}=\max\{\eta^{n_{0}},3\widetilde{c}_{0}\eta^{n_{0}}/2,2/c_{0},3\widetilde{c}_{0}/2,c/(m^{2}\delta_{0}),cm^{2}\}$, we complete the proof. ∎
@@
There exists a constant $C_{2}>1$ such that for any progressively overlapping sequence $\zeta\in(D_{b}^{m})^{\mathbb{N}_{0}}$ and for any totally distinct sequence $\xi\in(D_{b}^{m})^{\mathbb{N}_{0}}$, we have
$C_{2}^{-1}\lambda^{k}\leq|\Sigma_{k}^{\zeta}|\leq C_{2}\lambda^{k}\quad\text{and}\quad C_{2}^{-1}\eta^{k}\leq|\Sigma_{k}^{\xi}|\leq C_{2}\eta^{k}\quad\forall k\in\mathbb{N},$
where $\lambda$ and $\eta$ are defined as in Theorem 1.4.
@@
where $\widetilde{\mathbf{v}}_{0}=(|\Sigma_{m}^{\xi}|,|\Sigma_{m-1}^{\xi}|,\ldots,|\Sigma_{1}^{\xi}|)$. ∎
@@
For $n\in\mathbb{N}$, let $\zeta\in(D_{b}^{m})^{\mathbb{N}_{0}}$ be progressively overlapping at position-$k$ for all $1\leq k\leq n$, and let $\xi\in(D_{b}^{m})^{\mathbb{N}_{0}}$ be totally distinct at position-$k$ for all $1\leq k\leq n$. Then by the proof of Proposition 3.5, there exists a constant $C_{2}>1$ such that